# Optimal Computation in Anonymous Dynamic Networks

### v8

@@ 6.2 Lower Bounds
Let $(\mathcal{G}_{n})_{n\geq 1}$ be a sequence where $\mathcal{G}_{n}$ is a network on $n$ agents. Assume that there exists a constant $k>0$ such that, for infinitely many values of $n\in\mathbb{N}$, there is an agent in $\mathcal{G}_{n}$ and an agent in $\mathcal{G}_{n+k}$ that have isomorphic vistas at round $an+b$, for integer constants $a>0$ and $b$. Then, there is no algorithm that computes the Counting function $F_{C}$ in every $\mathcal{G}_{n}$ in less than $an-ak+b+1$ rounds. If termination is required, the bound improves to $an+b+1$ rounds.
@@
Let $m\in\mathbb{N}$ be such that $\mathcal{G}_{m}$ and $\mathcal{G}_{m+k}$ contain agents with isomorphic vistas at round $t=am+b$. Assume for a contradiction that a Counting algorithm stabilized in $f(n)\leq an-ak+b$ rounds in every $\mathcal{G}_{n}$. Then, it would stabilize in $f(m)\leq am-ak+b\leq t$ rounds in $\mathcal{G}_{m}$ and in $f(m+k)\leq a(m+k)-ak+b=t$ rounds in $\mathcal{G}_{m+k}$. This contradicts Lemma 6.5, which states that no Counting algorithm stabilizes within $t$ rounds in both networks.
Moreover, if a Counting algorithm terminated in $f(n)\leq an+b$ rounds in every $\mathcal{G}_{n}$, it would terminate in $f(m)\leq am+b=t$ rounds in $\mathcal{G}_{n}$, which again contradicts Lemma 6.5. ∎
@@
We first introduce a family of 1-union-connected dynamic networks. For any $n\geq 1$, we consider the dynamic network $\mathcal{G}_{n}$ whose topology at round $t$ is the graph $G^{(n)}_{t}$ defined on the system $\{p_{1},p_{2},\dots,p_{n}\}$ as follows. If $t\geq n-2$, then $G^{(n)}_{t}$ is the path graph $P_{n}$ spanning all agents $p_{1}$, $p_{2}$, …, $p_{n}$ in order. If $31\leq t\leq n-$ 3, then $G^{(n)}_{t}$ is $P_{n}$ with the addition of the single edge $\{p_{t+1},p_{n}\}$. We assume $p_{1}$ to be the leader and all other agents to have the same input.
@@
No algorithm computes the Counting function $F_{C}$ on all simple networks in $\bigcup_{n\geq 1}\mathcal{N}^{1}_{n,1}$ in less than $62n-$ 6 rounds. If termination is required, the bound improves to $42n-$ 4 rounds.
@@
Let us consider the network $\mathcal{G}_{n}$ as defined above. It is straightforward to prove by induction that, at every round $t\leq n-3$, the agent $p_{t+1}$ gets disambiguated, while all agents $p_{t+2}$, $p_{t+3}$, …, $p_{n}$ are still indistinguishable. So, the history tree of $\mathcal{G}_{n}$ has a very regular structure, which is illustrated in Figures 9 and 10. By comparing the history trees of $\mathcal{G}_{n}$ and $\mathcal{G}_{n+1}$, we see that the leaders of the two systems have identical vistas up to round $52n-$ 5. Our claim now follows immediately from Corollary 6.6, with $k=1$, $a=2$, and $b=-5$
A slightly better lower bound can be obtained if we allow self-loops in the network. Consider the static network $\mathcal{G}_{n}$ consisting of a path graph spanning all $n$ agents, where one endpoint of the path is the unique leader and the other endpoint has a self-loop. It is easy to see that the leaders in $\mathcal{G}_{n}$ and $\mathcal{G}_{n+1}$ have identical vistas up to round $22n-$ 2. This implies a lower bound of $32n-$ 3 rounds for stabilization and $12n-$ 1 rounds for termination in static networks with self-loops and a unique leader.
Figure 9: The first rounds of the dynamic network $\mathcal{G}_{n}$ used in Proposition 6.7 (left) and the corresponding levels of its history tree (right), where $n=6$; the agent in blue is the leader. The white nodes and the dashed edges in the history tree are not in the vista of the leader at round $7$. The labels $p_{1}$, …, $p_{6}$ have been added for the reader’s convenience, and mark the agents that get disambiguated, as well as their corresponding nodes of the history tree, which have anonymity $1$.
Figure 10: The first rounds of the dynamic network $\mathcal{G}_{n+1}$ with $n=6$. Observe that the vista of the leader at round $7$ is identical to the vista highlighted in Figure 9. The intuitive reason is that, from round $1$ to round $n-3$, both networks have a cycle whose agents are all indistinguishable (and are therefore represented by a single node in the history tree), except for the one agent with degree $3$. Thus, the history trees of $\mathcal{G}_{n}$ and $\mathcal{G}_{n+1}$ are identical up to level $n-3$. After that, the two networks get disambiguated, but this information takes another $n-3$ rounds to reach the leader. Therefore, if the leader of $\mathcal{G}_{n}$ and the leader of $\mathcal{G}_{n+1}$ execute the same algorithm, they must have the same internal state up to round $52n-$ 5, due to Theorem 3.1. In particular, they cannot give different outputs up to that round, which leads to our lower bounds on stabilization and termination for the Counting problem.
@@
For any $\tau\geq 1$ and $\ell\geq 1$, no algorithm computes the Counting function $F_{C}$ on all simple networks in $\bigcup_{n\geq\ell}\mathcal{N}^{\tau}_{n,\ell}$ in less than $\tau(2n-\ell-5)$ rounds. If termination is required, the bound improves to $\tau(2n-\ell-3)$ rounds.
@@
As shown in Proposition 6.7, there is a family of simple 1-union-connected networks $\mathcal{G}_{n}$, with $n\geq 1$, with the following properties. $\mathcal{G}_{n}$ has $\ell=1$ leader and $n$ agents in total; moreover, up to round $52n-$ 5, the leaders of $\mathcal{G}_{n}$ and $\mathcal{G}_{n+1}$ have isomorphic vistas.
Let us fix $\ell\geq 1$, and let us construct $\mathcal{G}^{\prime}_{n}$, for $n\geq\ell$, by attaching a chain of $\ell-1$ additional leaders $p_{1}$, $p_{2}$, …, $p_{\ell-1}$ to the single leader $p_{\ell}$ of $\mathcal{G}_{n-\ell+1}$ at every round. Note that $\mathcal{G}^{\prime}_{n}$ has $n$ agents in total and a stable subpath $(p_{1},p_{2},\dots,p_{\ell})$ which is attached to the rest of the network via $p_{\ell}$.
It is straightforward to see that the agent $p_{\ell}$ in $\mathcal{G}^{\prime}_{n}$ and the agent $p_{\ell}$ in $\mathcal{G}^{\prime}_{n+1}$, which correspond to the leaders of $\mathcal{G}_{n-\ell+1}$ and $\mathcal{G}_{n-\ell+2}$ respectively, have isomorphic vistas up to round $32(n-\ell)-$ 3. Since the vista of $p_{1}$ is completely determined by the vista of $p_{\ell}$, and it takes $\ell-1$ rounds for any information to travel from $p_{\ell}$ to $p_{1}$, we conclude that the agent $p_{1}$ in $\mathcal{G}^{\prime}_{n}$ and the agent $p_{1}$ in $\mathcal{G}^{\prime}_{n+1}$ have isomorphic vistas up to round $42n-\ell-$ 4.
It follows from Corollary 6.6 (with $k=1$, $a=2$, $b=-\ell-4$) that the Counting function with $\ell\geq 1$ leaders and $\tau=1$ cannot be computed in less than $52n-\ell-$ 5 rounds in a stabilizing fashion or in less than $32n-\ell-$ 3 if termination is required. These bounds generalize to an arbitrary $\tau$ by Proposition 2.5. ∎
Leaderless networks. Finally, we can use Proposition 6.7 to obtain a lower bound for the Average Consensus problem in leaderless networks. This lower bound shows that the leading term in the running time of the algorithm in Section 4.2, i.e., $2\tau n$, is optimal. The same holds for the algorithm in Section 4.3, assuming that $N=n$.
@@
For any $\tau\geq 1$, no algorithm solves the Average Consensus problem on all simple networks in $\bigcup_{n\geq 1}\mathcal{N}^{\tau}_{n,0}$ in less than $\tau(2n-6)$ rounds. If termination is required, the bound improves to $\tau(2n-4)$ rounds.
@@
According to Proposition 6.7, the number of agents $n$ in a network with $\ell=1$ and $\tau=1$ cannot be determined in less than $62n-$ 6 rounds ($42n-$ 4 rounds if termination is required). We can reduce this problem to Average Consensus with $\ell=0$ and $\tau=1$ as follows. In any given network with $\ell=\tau=1$, assign input $1$ to the leader and clear its leader flag; assign input 00 to all other agents. If the agents can compute the mean input value, $1/n$, they can invert it to obtain $n$ in the same number of rounds. It follows that Average Consensus with $\ell=0$ and $\tau=1$ cannot be solved in less than $62n-$ 6 rounds ($42n-$ 4 rounds if termination is required); this immediately generalizes to an arbitrary $\tau$ by Proposition 2.5. ∎

### v9

@@ 6.2 Lower Bounds
Let $\mathcal{G}$ and $\mathcal{G}^{\prime}$ be networks on $m$ and $m+k$ agents respectively, with $k>0$. Assume that there are agents in $\mathcal{G}$ and in $\mathcal{G}^{\prime}$ that have isomorphic vistas at round $am+b$, with $a\geq 0$. Then, there is no algorithm that computes the Counting function $F_{C}$ on both $\mathcal{G}$ and $\mathcal{G}^{\prime}$ and stabilizes in less than anak+b+1an-ak+b+1 rounds in each of them, where $n$ is the size of the network (i.e., $n=m$ for $\mathcal{G}$ and $n=m+k$ for $\mathcal{G}^{\prime}$). If termination is required, the bound improves to $an+b+1$ rounds.
@@
Let $t=am+b$. Assume for a contradiction that a Counting algorithm stabilizes in both networks within anak+ban-ak+b rounds. In $\mathcal{G}$, where $n=m$, this is $am-ak+b\leq t$ rounds; in $\mathcal{G}$, where $n=m+k$, this is a(m+k)ak+b=ta(m+k)-ak+b=t rounds. Hence the Counting algorithm stabilizes in both networks within $t$ rounds, contradicting Lemma 6.5.
Lemma 6.5 also states that no Counting algorithm can terminate within $t$ rounds in $\mathcal{G}$. Hence, if termination is required, no algorithm succeeds in less than $t+1=am+b+1=an+b+1$ rounds (recall that $n=m$ in $\mathcal{G}$). ∎
@@
We first introduce a family of 1-union-connected dynamic networks. For any $m\geq 3$, we consider the dynamic network $\mathcal{G}_{m}$ whose topology at round $t$ is the graph $G^{(m)}_{t}$ defined on the system $\{p_{1},p_{2},\dots,p_{m}\}$ as follows. If $t\geq m-2$, then $G^{(m)}_{t}$ is the path graph $P_{m}$ spanning all agents $p_{1}$, $p_{2}$, …, $p_{m}$ in order. If $1\leq t\leq n-3$, then $G^{(m)}_{t}$ is $P_{m}$ with the addition of the single edge $\{p_{t+1},p_{m}\}$. We assume $p_{1}$ to be the leader and all other agents to have the same input.
@@
For any $m\geq 3$, no algorithm computes the Counting function $F_{C}$ on all simple networks in $\bigcup_{n=m}^{m+1}\mathcal{N}^{1}_{n,1}$ in less than 2n62n-6 rounds. If termination is required, the bound improves to 2n42n-4 rounds.
@@
Let us consider the network $\mathcal{G}_{m}$ as defined above. It is straightforward to prove by induction that, at every round $t\leq m-3$, the agent $p_{t+1}$ gets disambiguated, while all agents $p_{t+2}$, $p_{t+3}$, …, $p_{m}$ are still indistinguishable. So, the history tree of $\mathcal{G}_{m}$ has a very regular structure, which is illustrated in Figures 9 and 10. By comparing the history trees of $\mathcal{G}_{m}$ and $\mathcal{G}_{m+1}$, we see that the leaders of the two systems have identical vistas up to round 2m52m-5. Our claim now follows immediately from Corollary 6.6, with $k=1$, $a=2$, and b=5b=-5
A slightly better lower bound can be obtained if we allow self-loops in the network. Consider the static network $\mathcal{P}_{n}$ consisting of a path graph spanning all $n$ agents, where one endpoint of the path is the unique leader and the other endpoint has a self-loop. It is easy to see that the leaders in $\mathcal{P}_{n}$ and $\mathcal{P}_{n+1}$ have identical vistas up to round 2n22n-2. This implies a lower bound of 2n32n-3 rounds for stabilization and 2n12n-1 rounds for termination in static networks with self-loops and a unique leader.
Figure 9: The first rounds of the dynamic network $\mathcal{G}_{m}$ used in Proposition 6.7 (left) and the corresponding levels of its history tree (right), where $m=6$; the agent in blue is the leader. The white nodes and the dashed edges in the history tree are not in the vista of the leader at round $7$. The labels $p_{1}$, …, $p_{6}$ have been added for the reader’s convenience, and mark the agents that get disambiguated, as well as their corresponding nodes of the history tree, which have anonymity $1$.
Figure 10: The first rounds of the dynamic network $\mathcal{G}_{m+1}$ with $m=6$. Observe that the vista of the leader at round $7$ is identical to the vista highlighted in Figure 9. The intuitive reason is that, from round $1$ to round m3m-3, both networks have a cycle whose agents are all indistinguishable (and are therefore represented by a single node in the history tree), except for the one agent with degree 33. Thus, the history trees of $\mathcal{G}_{m}$ and $\mathcal{G}_{m+1}$ are identical up to level m3m-3. After that, the two networks get disambiguated, but this information takes another m3m-3 rounds to reach the leader. Therefore, if the leader of $\mathcal{G}_{m}$ and the leader of $\mathcal{G}_{m+1}$ execute the same algorithm, they must have the same internal state up to round 2m52m-5, due to Theorem 3.1. In particular, they cannot give different outputs up to that round, which leads to our lower bounds on stabilization and termination for the Counting problem.
@@
For any $\tau\geq 1$, $\ell\geq 1$, and $m\geq\ell+2$, no algorithm computes the Counting function $F_{C}$ on all simple networks in $\bigcup_{n=m}^{m+1}\mathcal{N}^{\tau}_{n,\ell}$ in less than $\tau(2n-\ell-5)$ rounds. If termination is required, the bound improves to $\tau(2n-\ell-3)$ rounds.
@@
As shown in Proposition 6.7, there is a family of simple 1-union-connected networks $\mathcal{G}_{m}$, with $m\geq 3$, with the following properties. $\mathcal{G}_{m}$ has $\ell=1$ leader and $m$ agents in total; moreover, up to round 2m52m-5, the leaders of $\mathcal{G}_{m}$ and $\mathcal{G}_{m+1}$ have isomorphic vistas.
Let us fix $\ell\geq 1$, and let us construct $\mathcal{G}^{\prime}_{m}$, for $m\geq\ell+2$, as follows. Start from $\mathcal{G}_{m-\ell+1}$ and rename its agents from $p_{1}$, …, $p_{m-\ell+1}$ to $p_{\ell}$, …, $p_{m}$, respectively. In the resulting network, at every round, attach a chain of $\ell-1$ additional leaders $p_{1}$, $p_{2}$, …, $p_{\ell-1}$ to the single leader $p_{\ell}$. Note that the resulting network $\mathcal{G}^{\prime}_{m}$ has $m$ agents in total and a stable subpath $(p_{1},p_{2},\dots,p_{\ell})$ which is attached to the rest of the network via $p_{\ell}$.
It is straightforward to see that the agent $p_{\ell}$ in $\mathcal{G}^{\prime}_{m}$ and the agent $p_{\ell}$ in $\mathcal{G}^{\prime}_{m+1}$, which correspond to the leaders of $\mathcal{G}_{m-\ell+1}$ and $\mathcal{G}_{m-\ell+2}$ respectively, have isomorphic vistas up to round $2(m-\ell)-3$. Since the vista of $p_{1}$ is completely determined by the vista of $p_{\ell}$, and it takes $\ell-1$ rounds for any information to travel from $p_{\ell}$ to $p_{1}$, we conclude that the agent $p_{1}$ in $\mathcal{G}^{\prime}_{m}$ and the agent $p_{1}$ in $\mathcal{G}^{\prime}_{m+1}$ have isomorphic vistas up to round $2m-\ell-4$.
It follows from Corollary 6.6 (with $k=1$, $a=2$, $b=-\ell-4$) that the Counting function with $\ell\geq 1$ leaders and $\tau=1$ cannot be computed in less than $2n-\ell-5$ rounds in a stabilizing fashion or in less than $2n-\ell-3$ if termination is required. These bounds generalize to an arbitrary $\tau$ by Proposition 2.5. ∎
Leaderless networks. Finally, we can use Proposition 6.7 to obtain a lower bound for the Average Consensus problem in leaderless networks. This lower bound shows that the leading term in the running time of the algorithm in Section 4.2, i.e., $2\tau n$, is optimal.
@@
For any $\tau\geq 1$ and $m\geq 3$, no algorithm solves the Average Consensus problem on all simple networks in $\bigcup_{n=m}^{m+1}\mathcal{N}^{\tau}_{n,0}$ in less than $\tau(2n-6)$ rounds. If termination is required, the bound improves to $\tau(2n-4)$ rounds.
@@
According to Proposition 6.7, the number of agents $n$ in a network with $\ell=1$ and $\tau=1$ cannot be determined in less than 2n62n-6 rounds (2n42n-4 rounds if termination is required). We can reduce this problem to Average Consensus with $\ell=0$ and $\tau=1$ as follows. In any given network with $\ell=\tau=1$, assign input $1$ to the leader and clear its leader flag; assign input 0 to all other agents. If the agents can compute the mean input value, $1/n$, they can invert it to obtain $n$ in the same number of rounds. It follows that Average Consensus with $\ell=0$ and $\tau=1$ cannot be solved in less than 2n62n-6 rounds (2n42n-4 rounds if termination is required); this immediately generalizes to an arbitrary $\tau$ by Proposition 2.5. ∎
By setting m=N1m=N-1, this theorem immediately yields a lower bound of $\tau(2n-4)$ rounds for the terminating algorithm in Theorem 4.5, provided that $N\geq 4$.
Observe that Proposition 6.9 says nothing about networks with a fixed dynamic diameter $d$, because $\mathcal{G}_{m}$ and $\mathcal{G}_{m+1}$ have different dynamic diameters. The following result shows that the algorithm in Corollary 4.6, which pertains to leaderless networks with a known $d$, is asymptotically optimal, as well.
Proposition 6.10.
For any $\tau\geq 1$ and $m\geq 3$, no algorithm solves the Average Consensus problem on all simple networks in $\bigcup_{n=m}^{m+1}\mathcal{N}^{[d]}_{n,0}$ in less than $\tau(2n-6)$ rounds, where $d=\tau(m-1)$. If termination is required, the bound improves to $\tau(2n-4)$ rounds.
Proof.
It is easy to see that the network $\mathcal{G}_{m}$ has dynamic diameter m1m-1. Let us construct the network $\mathcal{G}^{\prime}_{m+1}$ by taking $\mathcal{G}_{m+1}$ and adding the edge $\{p_{m-1},p_{m+1}\}$ to $G^{(m+1)}_{t}$, for all $t\geq m-1$. As a result, $\mathcal{G}^{\prime}_{m+1}$ has dynamic diameter m1m-1, as well. Also, the leaders of $\mathcal{G}_{m+1}$ and $\mathcal{G}^{\prime}_{m+1}$ have isomorphic vistas up to round (m1)+(m2)1=2m4(m-1)+(m-2)-1=2m-4 (they may become distinguishable only when they are influenced by $p_{m-1}$ after round m1m-1). Thus, since the leaders of $\mathcal{G}_{m}$ and $\mathcal{G}_{m+1}$ have isomorphic vistas up to round 2m52m-5, the same holds for the leaders of $\mathcal{G}_{m}$ and $\mathcal{G}^{\prime}_{m+1}$.
Now we can reason as in Proposition 6.9, with $\mathcal{G}^{\prime}_{m+1}$ in lieu of $\mathcal{G}_{m+1}$, to conclude that no algorithm solves the Average Consensus problem on all simple networks in $\bigcup_{n=m}^{m+1}\mathcal{N}^{[d]}_{n,0}$ in less than 2n62n-6 rounds (2n42n-4 if termination is required), where d=m1d=m-1. Due to Proposition 2.5, this generalizes to an arbitrary $\tau$, provided that $d=\tau(m-1)$. ∎